# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

### v6

@@ 1 Introduction
Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9]. However, recent research has revealed that these models are vulnerable to typographic attacks [10, 11, 12], which exploit the reliance on the textual component of multimodal learning.
@@
With the growing use of VLMs in real-world applications, such as medical applications [13, 14, 15] and autonomous systems [16], as well as in Large Vision-Language Models (LVLMs) [17, 18, 19], it is essential to ensure the safety and reliability of these models.
@@
    LVLMs, such as the LLaVA family, exhibit vulnerabilities to typographic attacks, reflecting weaknesses inherited from their vision encoders.
@@
Table 1: Performance of VLMs and LVLMs available through OpenCLIP [20] resp. ollama [21], and Anthropic’s resp. OpenAI’s API on the SCAM datasets. Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding. Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model [22]. This table presents a only selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
@@ 2 Related Work
Multimodal pretraining, as established by CLIP [1] and SigLIP [2], leverages image-text pairs to capture rich semantics suitable for a variety of applications [23, 24]. This approach commonly features a contrastive objective to align visual and textual content in a shared latent space. Following this paradigm, various Vision-Language Models (VLMs) [25, 26, 27] have been proposed, demonstrating impressive performance in numerous downstream tasks [5, 6, 7]. Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [28, 29, 30, 31, 32, 33, 34, 35].
Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation [8, 9].
@@
While vision-language pretraining enables strong performance, research has shown that VLMs and LVLMs are susceptible to typographic attacks [10, 11, 12]. Existing works demonstrate several approaches to employ typographic attacks, for instance by Visual-Prompt-Injection [36], with self-generated attacks [12], or in scene-coherent fashion [37]. Since VLMs are integrated into autonomous systems, typographic attacks are also known to occur in safety-critical applications such as autonomous driving [38].
@@
To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works [39, 40, 11, 12, 36, 41] synthetically add attack words into existing image datasets while only few [42, 43, 44] propose datasets that exhibit real-world typographic attacks. Existing real-world datasets, such as PAINT [42] and Materzynska+[43], lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models. The previously largest real-world typographic attack dataset, RTA-100 [44], consists of 1000 images spanning 100 objects and attack words. However, these existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see LABEL:tab:dataset_comparison for a comparison). Our dataset addresses this limitation by introducing substantially larger variability across object–word combinations, which mitigates dataset bias and enables a more reliable assessment of generalization in VLMs and LVLMs. Furthermore, current datasets lack a structured comparison between real-world and synthetic typographic attacks, preventing a systematic analysis of transferability between these domains. SCAM is the first dataset to provide three explicitly aligned components: (i) SCAM, containing real-world attacked scenes, (ii) NoSCAM, the corresponding clean counterparts of the exact same scenes, and (iii) SynthSCAM, synthetically manipulated versions of the same object–word pairs. This tri-part design enables precise, controlled measurement of typographic attack effects and systematic comparisons across clean, real, and synthetic domains, establishing a stronger foundation for rigorous evaluation of VLMs and LVLMs as well as the development of effective defense mechanisms.
@@ 3 SCAM Dataset
To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images. This provides clean, paired counterparts for every attacked image. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images. Choosing a uniform font style in this synthetic variant was a deliberate simplification. Prior synthetic attack datasets have already explored variations in font position, opacity, and color [40], whereas our focus is on providing clean, attacked, and synthetic counterparts of the same scenes.
@@
All images were manually annotated with object and attack word labels as well as post-it size (in %) as part of a voluntary community effort. The annotation process was conducted using Label Studio [45]. To standardize the dataset, all images were zero-padded and downscaled to $512\times 512$ pixels using cv2.INTER_AREA [46], an interpolation method that preserves details while minimizing distortion.
@@
SCAM comprises 1162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations. The complete lists of objects and attack words are provided in section˜A.8. We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems. The distribution of these groups is depicted in fig.˜3.
@@
PAINT [42] 110 89 87
Materzynska+ [43] 180 20 8
RTA-100 [44] 1000 100 100
@@
As summarized in LABEL:tab:dataset_comparison, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words. Compared to RTA-100 [44], SCAM covers a much broader range of objects (660 vs. 100) and attack words (206 vs. 100), offering a more comprehensive evaluation of model robustness to typographic attacks. The three explicitly aligned components – SCAM, NoSCAM, and SynthSCAM – allow precise, controlled measurement of typographic attack effects and direct comparisons between clean, real, and synthetic conditions, thereby closing critical gaps in prior datasets.
@@ 4 Evaluation
We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup. Both evaluations are conducted on the SCAM datasets (SCAM, NoSCAM, SynthSCAM) and additionally on PAINT [42] and RTA-100 [44]. The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
@@
To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [47, 48, 1, 49, 20]. We also evaluate eleven LVLMs, including Claude Sonnet 4 [35], GPT-4o [28], LLaVA [29, 30, 31], Gemma3 [32], Llama3.2-vision [33] and Llama4 [34] models. The LLaVA, Gemma3, and Llama models are deployed via the Ollama platform [21], whereas Claude Sonnet 4 and GPT-4o is accessed via OpenAI’s resp. Anthropic’s API. An overview of selected models is provided in table˜1, while LABEL:tab:full-overview presents the complete list of all 111 evaluated models.
@@
We evaluate the performance of VLMs in a zero-shot classification task. For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word. These text embeddings are generated using the set of text templates proposed in [1], as detailed in section˜A.2. The predicted label is then determined based on the highest cosine similarity score. An overview of this process is depicted in fig.˜1b. Here, when the model is presented with an image of a “clock” with the attack word “taxi”, it does not classify the image correctly. All evaluations of VLMs were performed on an A100 40GB card.
@@
To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications. Our evaluation follows the setup illustrated in fig.˜1c, inspired by [50]. The model is provided with an image along with the following prompt:
@@
All LLaVA and Gemma3 models were evaluated on an A100 40GB card, and llama3.2-vision:90b and llama4:scout on an H100 80GB.
@@ 5 Results
We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table˜1. Our results show that typographic attacks on VLMs lead to an average drop in accuracy of 26 percentage points. LVLMs exhibit similar behavior, with the choice of vision encoder appearing to influence the extent of accuracy degradation. Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.˜4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table˜4. These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
@@
((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
SCAM fails $648\pm 192$ $154\pm 68$
SCAM succeeds $18\pm 11$ $341\pm 151$
((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
SCAM fails $741\pm 243$ $96\pm 47$
SCAM succeeds $124\pm 71$ $201\pm 137$
@@
((a)) Average across different CLIP and SigLIP models trained by OpenAI and Visheratin [51] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1). More details about the selection process can be found in section˜A.4.
@@
Table˜3 presents the impact of training data on the ViT-B models [52] trained by the OpenCLIP team. We focus solely on these models as they share the same training paradigm but are trained on different datasets. Overall, the ViT-B models trained on the LAION datasets [47] are highly susceptible to typographic attacks, while certain CommonPool variants [53] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively. Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI. Consequently, as indicated by the accuracies for that model in table˜1, this filtering criterion appears to propagate susceptibility.
While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance. We compare different architectures: CLIP-RN (ResNets [54]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [51] (see table˜5(a)). Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models. Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
Lastly, fig.˜6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks. However, we find that the original CLIP models [1] trained by OpenAI perform significantly worse across all sizes.
@@
We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig.˜6). The mean accuracy across all models follows this trend, which is consistent with the findings in [40].
@@
Section˜A.6 presents a preliminary experiment that isolates the role of the vision encoder in LVLMs with respect to typographic robustness. The experiment does not show a clear gain in robustness when substituting the vision encoder with one ranked safer by SCAM, while keeping the LLM, training data, and training regime unchanged. Instead, the overall performance decreases, highlighting the difficulty of choosing the optimal vision encoder and showcasing the need for further experiments investigating the robustness impact of different VLM choices. Apart from that, as discussed in section˜A.7, scaling up the LLM backbone not only improves performance on SCAM but also enhances the model’s general typographic understanding on textual and OCR-based benchmarks.

### v7

@@ Abstract
Keywords: Typographic Attack Dataset, Vision-Language Models, Benchmarking
@@ 1 Introduction
Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) have demonstrated remarkable performance across various multimodal tasks (Gan et al., 2022), enabling applications in image classification (Radford et al., 2021; Menon and Vondrick, 2022), multimodal retrieval (Bogolin et al., 2022; Ming and Li, 2024), and generative tasks (Ramesh et al., 2022; Rombach et al., 2022; Gafni et al., 2022). However, recent research has revealed that these models are vulnerable to typographic attacks (Goh et al., 2021; Cheng et al., 2024c; Qraitem et al., 2024), which exploit the reliance on the textual component of multimodal learning.
@@
With the growing use of VLMs in real-world applications, such as medical applications (Zhao et al., 2023; Li et al., 2023b; Hartsock and Rasool, 2024) and autonomous systems (Zhou et al., 2024), as well as in Large Vision-Language Models (LVLMs) (Liu et al., 2024c, a; Dubey et al., 2024), it is essential to ensure the safety and reliability of these models.
@@
    LVLMs, such as the LLaVA family, exhibit vulnerabilities to typographic attacks, reflecting weaknesses present in their vision encoders.
@@
Table 1: Performance of VLMs and LVLMs available through OpenCLIP (OpenCLIP, ) resp. ollama (Ollama, ), and Anthropic’s resp. OpenAI’s API on the SCAM datasets. Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding. Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model (Razzhigaev et al., 2023). This table presents a only selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
@@ 2 Related Work
Multimodal pretraining, as established by CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), leverages image-text pairs to capture rich semantics suitable for a variety of applications (Dorbala et al., 2022; Zhou et al., 2023). This approach commonly features a contrastive objective to align visual and textual content in a shared latent space. Following this paradigm, various Vision-Language Models (VLMs) (Jia et al., 2021; Li et al., 2022; Tschannen et al., 2025) have been proposed, demonstrating impressive performance in numerous downstream tasks (Bogolin et al., 2022; Ramesh et al., 2022; Ming and Li, 2024). Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks (Hurst et al., 2024; Liu et al., 2023b, a, 2024b; Team et al., 2025; Grattafiori et al., 2024; Meta, 2025; Anthropic, 2025).
Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation (Rombach et al., 2022; Gafni et al., 2022).
@@
While vision-language pretraining enables strong performance, research has shown that VLMs and LVLMs are susceptible to typographic attacks (Goh et al., 2021; Cheng et al., 2024c; Qraitem et al., 2024). Existing works demonstrate several approaches to employ typographic attacks, for instance by Visual-Prompt-Injection (Kimura et al., 2024), with self-generated attacks (Qraitem et al., 2024), or in scene-coherent fashion (Cao et al., 2024b). Since VLMs are integrated into autonomous systems, typographic attacks are also known to occur in safety-critical applications such as autonomous driving (Chung et al., 2024).
@@
To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works (Gong et al., 2023; Cheng et al., 2024d, a, c; Qraitem et al., 2024; Kimura et al., 2024) synthetically add attack words into existing image datasets while only few (Ilharco et al., 2022; Materzyńska et al., 2022; Azuma and Matsui, 2023) propose datasets that exhibit real-world typographic attacks. Existing real-world datasets, such as PAINT (Ilharco et al., 2022) and Materzynska+(Materzyńska et al., 2022), lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models. The previously largest real-world typographic attack dataset, RTA-100 (Azuma and Matsui, 2023), consists of 1000 images spanning 100 objects and attack words. However, these existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see LABEL:tab:dataset_comparison for a comparison, and section˜A.2 for example images of PAINT and RTA-100). Our dataset addresses this limitation by introducing substantially larger variability across object–word combinations, which mitigates dataset bias and enables a more reliable assessment of generalization in VLMs and LVLMs. Furthermore, current datasets lack a structured comparison between real-world and synthetic typographic attacks, preventing a systematic analysis of transferability between these domains. SCAM is the first dataset to provide three explicitly aligned components: (i) SCAM, containing real-world attacked scenes, (ii) NoSCAM, the corresponding clean counterparts of the exact same scenes, and (iii) SynthSCAM, synthetically manipulated versions of the same object–word pairs. This tri-part design enables precise, controlled measurement of typographic attack effects and systematic comparisons across clean, real, and synthetic domains, establishing a stronger foundation for rigorous evaluation of VLMs and LVLMs as well as the development of effective defense mechanisms.
@@ 3 SCAM Dataset
To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images. This provides clean, paired counterparts for every attacked image. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images. Choosing a uniform font style in this synthetic variant was a deliberate simplification. Prior synthetic attack datasets have already explored variations in font position, opacity, and color (Cheng et al., 2024a), whereas our focus is on providing clean, attacked, and synthetic counterparts of the same scenes.
@@
All images were manually annotated with object and attack word labels as well as post-it size (in %) as part of a voluntary community effort. The annotation process was conducted using Label Studio (Studio, ). To standardize the dataset, all images were zero-padded and downscaled to $512\times 512$ pixels using cv2.INTER_AREA (OpenCV, ), an interpolation method that preserves details while minimizing distortion.
@@
SCAM comprises 1162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations. The complete lists of objects and attack words are provided in section˜A.9. We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems. The distribution of these groups is depicted in fig.˜3.
@@
PAINT (Ilharco et al., 2022) 110 89 87
Materzynska+ (Materzyńska et al., 2022) 180 20 8
RTA-100 (Azuma and Matsui, 2023) 1000 100 100
@@
As summarized in LABEL:tab:dataset_comparison, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words. Compared to RTA-100 (Azuma and Matsui, 2023), SCAM covers a much broader range of objects (660 vs. 100) and attack words (206 vs. 100), offering a more comprehensive evaluation of model robustness to typographic attacks. The three explicitly aligned components – SCAM, NoSCAM, and SynthSCAM – allow precise, controlled measurement of typographic attack effects and direct comparisons between clean, real, and synthetic conditions, thereby closing critical gaps in prior datasets.
Maintenance and Ethical Use.
SCAM is publicly hosted on Hugging Face under a stable organization account [^2] https://huggingface.co/datasets/BLISS-e-V/SCAM. The dataset card provides versioning, update logs, and contact information, and we will maintain SCAM by releasing documented updates if errors are discovered or extensions are added. Users can contact the corresponding authors to report issues, request corrections, or raise takedown concerns.
All images in SCAM depict everyday objects with handwritten English words on post-it notes. SCAM does not contain personal, biometric, or otherwise sensitive data. The dataset is released for research purposes under the license specified on the dataset card, and users are expected to follow standard citation practices and refrain from deploying models trained on SCAM in applications intended to harm or deceive humans. A more detailed discussion of potential risks and benefits is provided in the Broader Impact Statement on page Broader Impact Statement.
@@ 4 Evaluation
We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup. Both evaluations are conducted on the SCAM datasets (SCAM, NoSCAM, SynthSCAM) and additionally on PAINT (Ilharco et al., 2022) and RTA-100 (Azuma and Matsui, 2023). The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
@@
To evaluate VLMs such as CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), we use the OpenCLIP suite (Ilharco et al., 2021; Schuhmann et al., 2022; Radford et al., 2021; Cherti et al., 2023; OpenCLIP, ). We also evaluate eleven LVLMs, including Claude Sonnet 4 (Anthropic, 2025), GPT-4o (Hurst et al., 2024), LLaVA (Liu et al., 2023b, a, 2024b), Gemma3 (Team et al., 2025), Llama3.2-vision (Grattafiori et al., 2024) and Llama4 (Meta, 2025) models. The LLaVA, Gemma3, and Llama models are deployed via the Ollama platform (Ollama, ), whereas Claude Sonnet 4 and GPT-4o is accessed via OpenAI’s resp. Anthropic’s API. An overview of selected models is provided in table˜1, while LABEL:tab:full-overview presents the complete list of all 111 evaluated models.
@@
We evaluate the performance of VLMs in a zero-shot classification task. For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word. These text embeddings are generated using the set of text templates proposed in (Radford et al., 2021), as detailed in section˜A.3. The predicted label is then determined based on the highest cosine similarity score. An overview of this process is depicted in fig.˜1b. Here, when the model is presented with an image of a “clock” with the attack word “taxi”, it does not classify the image correctly. All evaluations of VLMs were performed on an A100 40GB card.
@@
To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications. Our evaluation follows the setup illustrated in fig.˜1c, inspired by (Cheng et al., 2024b). The model is provided with an image along with the following prompt:
@@
All LLaVA and Gemma3 models were evaluated on an A100 40GB card, and
llama3.2-vision:90b and llama4:scout on an H100 80GB.
((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
SCAM fails $648\pm 192$ $154\pm 68$
SCAM succeeds $18\pm 11$ $341\pm 151$
((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
SCAM fails $741\pm 243$ $96\pm 47$
SCAM succeeds $124\pm 71$ $201\pm 137$
@@ 5 Results
We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table˜1. Our results show that typographic attacks on VLMs lead to an average drop in accuracy of 26 percentage points. LVLMs exhibit similar behavior, with the choice of vision encoder appearing to influence the extent of accuracy degradation. Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.˜4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table˜3. These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
@@
Table˜4 presents the impact of training data on the ViT-B models (Dosovitskiy et al., 2020) trained by the OpenCLIP team. We focus solely on these models as they share the same training paradigm but are trained on different datasets. Overall, the ViT-B models trained on the LAION datasets (Schuhmann et al., 2022) are highly susceptible to typographic attacks, while certain CommonPool variants (Gadre et al., 2023) exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively. Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 (Radford et al., 2021) model from OpenAI. Consequently, as indicated by the accuracies for that model in table˜1, this filtering criterion appears to propagate susceptibility.
While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance. We compare different architectures: CLIP-RN (ResNets (He et al., 2016)) and CLIP-ViT models trained by OpenAI (Radford et al., 2021), as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB (Visheratin, 2023) (see table˜5(a)). Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models. Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
@@
((a)) Average across different CLIP and SigLIP models trained by OpenAI and Visheratin (2023) on LAION-COCO NLLB (referred to, in OpenCLIP, as v1). More details about the selection process can be found in section˜A.5.
@@
Lastly, fig.˜6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks. However, we find that the original CLIP models (Radford et al., 2021) trained by OpenAI perform significantly worse across all sizes.
@@
We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig.˜6). The mean accuracy across all models follows this trend, which is consistent with the findings in (Cheng et al., 2024a).
@@
Section˜A.7 presents a preliminary experiment that isolates the role of the vision encoder in LVLMs with respect to typographic robustness. The experiment does not show a clear gain in robustness when substituting the vision encoder with one ranked safer by SCAM, while keeping the LLM, training data, and training regime unchanged. Instead, the overall performance decreases, highlighting the difficulty of choosing the optimal vision encoder and showcasing the need for further experiments investigating the robustness impact of different VLM choices. Apart from that, as discussed in section˜A.8, scaling up the LLM backbone not only improves performance on SCAM but also enhances the model’s general typographic understanding on textual and OCR-based benchmarks.